A 100.00 mL sample of spring water was treated to


Help!
Answer #1

We have to convert the 25.00mL of 0.002107M K2Cr2O7 to moles (remember to use volume in liters):

total moles = M*L = (0.002107 mol/L)(0.025L)

total moles = 0.000052675 moles K2Cr2O7

Those are the total moles of K2Cr2O7 used in the titration, however there was an excess which was titrated with Fe2+, we have to calcultate the moles of K2Cr2O7 in excess so we can know the moles of K2Cr2O7 which reacted with iron in the water sample:

We calculate the Fe2+ moles that reacted with the excess K2Cr2O7 :

moles Fe2+ = M*L

moles Fe2+ = (0.00979 mol/L) (0.00747L)

moles Fe2+ = 0.0000731313 moles Fe2+

Now, using the coefficients in the balanced chemical equation, we calculate the excess moles of K2Cr2O7

excess moles K2Cr2O7 = 0.0000731313 moles Fe2+ (1 mol K2Cr2O7 / 6 mol Fe2+)

excess moles K2Cr2O7 = 0.00001218855 moles K2Cr2O7

Now we calculate the moles of K2Cr2O7 that reacted with iron in the water, it's just a substraction:

total moles K2Cr2O7 - excess moles K2Cr2O7 = reaction moles K2Cr2O7

reaction moles K2Cr2O7 = 0.00004048645 moles K2Cr2O7

We convert those moles to moles of Fe2+ present in the sample, again using the coefficients in the balanced equation:

moles Fe2+ (sample) = 0.00004048645 moles K2Cr2O7 (6 mol Fe2+ / 1mol  K2Cr2O7)

moles Fe2+ (sample) = 0.0002429187 moles Fe2+

To calculate ppm we convert those moles to miligrams (using molar mass of Fe = 55.845g/mol) and multiplying by 1000 (to convert to miligrams) and divide by sample volume (0.1L)

miligrams Fe2+ = 0.0002429187 moles Fe2+ (55.845g / mol) (1000mg/1g)

miligrams Fe2+ = 13.566 mg

ppm = mg/L

ppm = 13.6mg/0.1L

ppm = 136ppm Fe2+